# Quantum automorphism groups of 2-graphs

### v2

@@ Abstract.
We formulate a notion of the quantum automorphism groups of $2$-graphs. We show that two isomorphic $2$-graphs have isomorphic quantum automorphism groups making the notion well-defined. We compute the quantum automorphism groups of a few $2$-graphs. It is observed that quantum automorphism group of a $1$-graph can be realized as the quantum automorphism group of a $2$-graph.
@@
higher rank graphs, quantum automorphism groups, quantum groups, $C^{*}$-algebra
@@ 1. Introduction
Higher rank graphs (synonymously rank $k$ graphs or $k$-graphs) are natural generalization of graphs. Roughly speaking, a $k$-graph $\Lambda$ is a countable small category equipped with a “degree map” $d:\Lambda\rightarrow\mathbb{N}^{k}$ with the morphisms admitting a unique factorization property. Since its inception, it has played a key role in the realm of noncommutative geometry and topology. In particular, it helps in the proper generalization of Cuntz-Krieger $C^{\ast}$-algebras to higher-rank graph $C^{\ast}$-algebras. There is a formidable literature on higher rank graphs and their $C^{\ast}$-algebras (see [NYJM], [Steger1], [Steger2], [Sims], [Sims1], [Sims2], [Hazzlewood1] to name a few). Pioneered by S. Wang ([Wang]), the quantum symmetry of various finite dimensional or infinite dimensional classical or quantum objects has been quite extensively studied by many authors in recent years (e.g. [Goswami], [Banica], [Hajac], [Skalski], [Asfaq], [Voigt2] [Vaes]). Such quantum symmetry is now well established as the generalized symmetry in the world of noncommutative geometry or topology. Among the finite structures, the quantum automorphism group of finite graphs (see [Schmidt], [Fulton]), at least for small number of vertices, are reasonably well understood by now. Moreover, the notion of the quantum automorphism groups of graphs has found its way into quantum information theory (see [Qinfo]) as well as some concrete physical models involving graphs (see [Potts]). Therefore, higher rank graphs being a generalization of usual graphs, it is quite natural to consider the notion of quantum automorphism groups of higher rank graphs. Recall that the quantum automorphism groups of $C^{\ast}$-algebras are typically defined to be the universal objects of certain category of compact quantum groups acting faithfully on the $C^{\ast}$-algebras. Existence of the universal objects in such categories are often tricky and requires careful consideration. In this paper, we define the notion of “the” quantum automorphism group of rank $2$ graphs. We restrict our attention to the class of $2$-graphs as they have an equivalent formulation in terms of a simpler combinatorial data circumventing the category theoretic machinery. Briefly speaking, a $2$-graph can be described by a triple consisting of two $1$-graphs on the same vertex set having commuting vertex matrices and a bijection between certain sets. We would like to mention that we restrict our attention to the case where the common vertex set has finitely many points and the individual graphs are without multiple edge, source or sink to be able to work in the category of compact quantum groups. Then the quantum automorphism group of such a triple is naturally defined. But one has to be careful as the isomorphism class of $2$-graphs depend on a natural notion of equivalence between triples. We consider the right notion of two equivalent triples and prove that the quantum automorphism groups of such equivalent triples are isomorphic. This allows us to call the quantum automorphism group of a triple “the” quantum automorphism group of the corresponding $2$-graph. We then compute the quantum automorphism group of some $2$-graphs. It is clear from the definition that in order to find the quantum automorphism group of a $2$-graph to be genuine one has to at least start with a triple which has two constituent graphs admitting genuine quantum automorphisms. But that is naturally not sufficient though as seen in Section 4 Example (b). A particular class of $2$-graphs are of importance where the constituent graphs of the triple are two copies of the same graph. In such cases if the bijection is the trivial one then obviously one gets the quantum automorphism group of the $1$-graph as the quantum automorphism group of the corresponding $2$-graph. In that way one can easily ensure genuine quantum symmetry by choosing the $1$-graph with genuine quantum symmetry. However, the case of non-trivial bijection is clearly more interesting. We have an example with a non-trivial bijection where the quantum automorphism group of the resulting $2$-graph is still genuine. Also we exhibit one example where we have two copies of the complete graph on $4$-vertices as the constituent graphs of the triple, but there is a choice of a bijection so that the quantum automorphism group of the corresponding triple is not only classical but also it is trivial. We end this paper with a brief discussion on possible problems to explore in future. We have also added an Appendix section where a brief sketch of the proof of the fact that the isomorphism classes of $2$-graphs depend on an equivalence class of triples has been included.
@@ 2.1. Higher rank graphs
Definition 2.2.
(see Definition $1.1$ of [NYJM]) A $k$-graph (rank $k$ graph or higher rank graph) $(\Lambda,d)$ consists of a countable small category $\Lambda$ (with range and source maps $r$ and $s$ respectively) together with a functor $d:\Lambda\rightarrow\mathbb{N}^{k}$ satisfying the factorisation property: for every morphism $\lambda\in\Lambda$ and $m,n\in\mathbb{N}^{k}$ with $d(\lambda)=m+n,$ there are unique elements $\mu,\nu\in\Lambda$ such that $\lambda=\mu\nu$ and $d(\mu)=m,\ d(\nu)=n.$
@@
Definition 2.4.
(see Definition $1.9$ of [NYJM]) Let $f:\mathbb{N}^{l}\rightarrow\mathbb{N}^{k}$ be a monoid morphism. If $(\Lambda,d)$ is a $k$-graph we may form the $l$-graph $f^{\ast}(\Lambda)$ as follows: $f^{\ast}(\Lambda)=\{(\lambda,n):d(\lambda)=f(n)\}\ {\rm with}\ d(\lambda,n)=n,s(\lambda,n)=s(\lambda)\ {\rm and}\ r(\lambda,n)=r(\lambda)$.
Now let us specialize to the case where $k=2$. We shall take an alternative equivalent view point of $2$-graphs which will be amenable to the study of quantum symmetry. This view point is described in Section $6$ of [NYJM]. Here we shall give the bare bones. Roughly speaking, any $2$-graph can be constructed from two $1$-graphs with common set of vertices and commuting vertex matrices. Again we remind the reader that in our case, since ${\rm obj}(\Lambda)$ is finite, the number of common vertices are finite.
$2$-graph: Let $\mathcal{G}_{1}=(V,E_{1})$, $\mathcal{G}_{2}=(V,E_{2})$ be two $1$-graphs with common set of vertices $V$ such that the corresponding vertex matrices and commute. The commutation of $A_{1}$ and $A_{2}$ guarantees the existence of a bijection $\theta:E_{1}\star E_{2}\rightarrow E_{2}\star E_{1}$ where
$E_{1}\star E_{2}=\{(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\times E_{2}|s(e^{1}_{\alpha})=t(e^{2}_{\beta})\},$
$E_{2}\star E_{1}=\{(e^{2}_{\alpha},e^{1}_{\beta})\in E_{2}\times E_{1}|s(e^{2}_{\alpha})=t(e^{1}_{\beta})\}.$
We call the data $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ a triple. Given such a triple, one can construct a $2$-graph $\Lambda$ in the following way:
The set ${\rm obj}(\Lambda)$ or $\Lambda^{(0,0)}$ is the set of common vertices of the graphs $\mathcal{G}_{1},\mathcal{G}_{2}$. The morphisms of degree $(1,0)$ and $(0,1)$ are determined by the edge sets $E_{1},E_{2}$. The higher degree morphisms are built step by step using the bijection $\theta$. For details of the construction, we again refer the readers to the Section 6 of [NYJM]. Conversely given a $2$-graph $\Lambda$, one can construct a triple. Roughly one takes the set $\Lambda^{(0,0)}$ to be the set of common vertices $V$. Then the morphisms $\Lambda^{(1,0)}$ and $\Lambda^{(0,1)}$ together with the range and source maps defines the constituent $1$-graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The bijection $\theta$ is then built out of the morphisms $\Lambda^{(1,1)}$ using the unique factorization property of the morphisms. A triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ corresponding to a $2$-graph $\Lambda$ will be called the defining triple of $\Lambda$.
@@
(1) A bijection between the sets $E_{1}\star E_{2}$ and $E_{2}\star E_{1}$ is guaranteed precisely because the vertex matrices commute. Also, $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$Rightarrow
$s(e^{1}_{\alpha})=t(e^{2}_{\beta});\ s(e^{2}_{\mu})=t(e^{1}_{\nu});$
$t(e^{1}_{\alpha})=t(e^{2}_{\mu});\ s(e^{2}_{\beta})=s(e^{1}_{\nu}).$
@@
(2) We also mention that we restrict our attention to the $2$-graphs where the individual $1$-graphs are without multiple edges, sinks or sources for technical reasons.
@@
Now let $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ be two triples. We denote the common vertex set of the triples by $V_{1},V_{2}$ respectively; the vertex matrices of $\mathcal{G}_{i}$ by $A_{i}$ and vertex matrices of $\mathcal{H}_{i}$ by $B_{i}$ for $i=1,2$ respectively; the edge sets of $\mathcal{G}_{i}$ and $\mathcal{H}_{i}$ by $E_{i}$ and $F_{i}$ for $i=1,2$ respectively.
@@
Two triples $(\mathcal{G}_{1},\mathcal{H}_{1},\theta_{1})$ and $(\mathcal{G}_{2},\mathcal{H}_{2},\theta_{2})$ as above are said to be equivalent if the following happens:
(i) There is a bijection $T:V_{1}\rightarrow V_{2}$. This means effectively we can assume that all the $1$-graphs involved share a common set of vertices $V$ and $T$ is a permutation of the vertices so that we have an invertible linear map $T_{\ast}:C(V)\rightarrow C(V)$.
(ii) If we denote the permutation matrix of the linear map $T_{\ast}$ again by $T$ with respect to standard bases of $C(V)$ then $TA_{2}=A_{1}T$ and $TB_{2}=B_{1}T$.
(iii) Consider the map $T^{\otimes 4}\equiv T_{\ast}\otimes T_{\ast}\otimes T_{\ast}\otimes T_{\ast}:C(V\times V\times V\times V)\rightarrow C(V\times V\times V\times V)$. Thanks to condition (ii), $T^{\otimes 4}$ is also a bijection between both $C(E_{1}\star F_{1})$, $C(E_{2}\star F_{2})$ and $C(F_{1}\star E_{1})$, $C(F_{2}\star E_{2})$. We have the following commutative diagram:
{tikzcd}
@@
We abbreviate the last condition of the above definition as $\theta$-compatibility. The isomorphism class of $2$-graphs depend only on the equivalence class of triples as defined above. We have included a brief sketch of the proof. We believe that it must be known. But since we could not find an explicit mention of this result in the literature, we have decided to keep the result in the Appendix (see Theorem A.1).
@@ 2.9. Compact quantum groups and their actions
The term quantum group is quite general. We make it clear that we work with Compact quantum groups (CQG in short) ala S.L. Woronowicz ([Woro]). The literature on CQG and their actions on $C^{\ast}$-algebras are aplenty by now. The reader can refer to the articles [DeCom], [Soltan] for details on the actions of CQG’s on $C^{\ast}$-algebras. Here we briefly recall what is relevant for us. The following are mainly taken from [edin].
@@
A compact quantum group (CQG) $\mathbb{G}$ is a pair $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ such that $C(\mathbb{G})$ is a unital $C^{\ast}$-algebra and $\Delta_{\mathbb{G}}:C(\mathbb{G})\rightarrow C(\mathbb{G})\otimes C(\mathbb{G})$ is a unital $C^{\ast}$- homomorphism satisfying
(i) $({\rm id}\otimes\Delta_{\mathbb{G}})\circ\Delta_{\mathbb{G}}=(\Delta_{\mathbb{G}}\otimes{\rm id})\circ\Delta_{\mathbb{G}}.$
(ii) ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(1\otimes C(\mathbb{G}))\}$ and ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(C(\mathbb{G})\otimes 1)\}$ are dense in $C(\mathbb{G})\otimes C(\mathbb{G})$.
@@
(i) $(\alpha\otimes{\rm id})\circ\alpha=({\rm id}\otimes\Delta_{\mathbb{G}})\circ\alpha$.
(ii) Span$\{\alpha({\cal C})(1\otimes\Delta_{\mathbb{G}})\}$ is dense in ${\cal C}\otimes C(\mathbb{G})$.
(iii) The $\ast$-algebra generated by the set $\{(\omega\otimes{\rm id})\circ\alpha({\cal C}):\omega\in{\cal C}^{\ast}\}$ is norm-dense in $C(\mathbb{G})$.
Definition 2.13.
(Def 2.1 of [Bichon]) Given a unital $C^{\ast}$-algebra ${\cal C}$, quantum automorphism group of ${\cal C}$ is a CQG $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ acting faithfully on ${\cal C}$ satisfying the following universal property:
@@
In general the universal object might fail to exist in the above category. To ensure the existence one generally assumes that the action preserves some fixed state on the $C^{\ast}$-algebra. However for certain finite dimensional $C^{\ast}$-algebras such a state preserving condition is automatic. We would not go into further details as we are not going to use it in this paper. For further details, the reader might consult [Wang].
@@
If we take the space of $n$ points $X_{n}$ then the quantum automorphism group of the $C^{\ast}$-algebra $C(X_{n})$ is given by the CQG (denoted by $S_{n}^{+}$) whose underlying $C^{\ast}$-algebra is the universal $C^{\ast}$ algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations (see Theorem 3.1 of [Wang]):
@@
Any $n\times n$ matrix $((q_{ij}))_{i,j=1,\ldots,n}$ with $q_{ij}$’s in a $C^{\ast}$-algebra satisfying the above relations is called a magic unitary. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. It is worth mentioning that the CQG $S_{n}^{+}$ is a particular type of CQG known as the compact matrix pseudogroup (see [Woro]). The $\ast$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ is the canonical dense Hopf $\ast$-algebra and the antipode is given on the generators by $\kappa(q_{ij})=q_{ji}$.

### v3

# Quantum isomorphism of 2-graphs

@@ Abstract.
We formulate a notion of the quantum automorphism group of a $2$-graph. After some preliminary computations, we define quantum isomorphism between a pair of $2$-graphs. We produce a ‘non-trivial’ example of a pair of $2$-graphs that are not quantum isomorphic to each other.
@@
higher rank graphs, quantum automorphism groups, quantum isomorphism, $C^{*}$-algebra
@@ 1. Introduction
Higher rank graphs (synonymously rank $k$ graphs or $k$-graphs) are natural generalization of graphs. Roughly speaking, a $k$-graph $\Lambda$ is a countable small category equipped with a “degree map” $d:\Lambda\rightarrow\mathbb{N}^{k}$ with the morphisms admitting a unique factorization property. Since its inception, it has played a key role in the realm of noncommutative geometry and topology. In particular, it helps in the proper generalization of Cuntz-Krieger $C^{\ast}$-algebras to higher-rank graph $C^{\ast}$-algebras. There is a formidable literature on higher rank graphs and their $C^{\ast}$-algebras (see [NYJM, Steger1, Steger2, Sims, Sims1, Sims2, Hazzlewood1] to name a few). Pioneered by S. Wang ([Wang]), the quantum symmetry of various finite dimensional or infinite dimensional classical or quantum objects has been quite extensively studied by many authors in recent years (e.g. [Goswami, Banica, Hajac, Skalski, Asfaq, Voigt2, Vaes]). Such quantum symmetry is now well established as the generalized symmetry in the world of noncommutative geometry or topology. Among the finite structures, the quantum automorphism group of finite graphs (see [Schmidt, Fulton]), at least for small number of vertices, are reasonably well understood by now. Moreover, the notion of the quantum automorphism groups of graphs has found its way into quantum information theory (see [Qinfo, Qinfo2]) as well as some concrete physical models involving graphs (see [Potts]). Therefore, higher rank graphs being a generalization of usual graphs, it is quite natural to consider the notion of quantum automorphism groups of higher rank graphs.
In this paper, we define the notion of “the” quantum automorphism group of a rank $2$ graph. The notion is, in a sense, quite natural. Then, with a growing list of literature in the intersection of quantum groups and nonlocal games, we are naturally led to define quantum isomorphism between a pair of $2$-graphs. We restrict our attention to the class of $2$-graphs as they have an equivalent formulation in terms of a simpler combinatorial data circumventing the category theoretic machinery. Briefly speaking, a $2$-graph can be described by a triple consisting of two $1$-graphs on the same vertex set having commuting vertex matrices and a bijection between certain sets. We would like to mention that we restrict our attention to the case where the common vertex set has finitely many vertices to be able to work in the category of compact quantum groups. Then the quantum automorphism group of such a triple is naturally defined. But one has to be careful as the isomorphism class of $2$-graphs depend on a natural notion of equivalence between triples. Then the quantum automorphism groups of such equivalent triples are isomorphic. This allows us to call the quantum automorphism group of a triple “the” quantum automorphism group of the corresponding $2$-graph.
Let us briefly mention the organization of the paper and some of the main results obtained here. We start with a preliminary section recalling the basics of higher rank (in particular rank $2$) graphs and compact quantum groups. Then in the third section, we define the notion of quantum automorphism group of a $2$-graph. The underlying $C^{*}$-algebra of the quantum automorphism group is described in terms of generators and relations (see Equation (3.3) for a new kind of relations). The next section (section 4) is devoted to some preliminary computations. It is clear from the definition that in order to ensure that the quantum automorphism group of a $2$-graph is genuine one has to at least start with a triple which has two constituent graphs admitting genuine quantum automorphisms. But that is naturally not sufficient though as seen in Example (b). It is also clear from the definition that if one starts with two copies of the same $1$-graph and the defining bijection as the trivial bijection, the quantum automorphism group becomes the usual quantum automorphism group of the constituent $1$-graph. Thus when one considers two copies of the same graph, a non-trivial bijection becomes interesting. We exhibit one example where we have two copies of the complete graph on $4$-vertices as the constituent graphs of the triple, but there is a choice of a bijection so that the quantum automorphism group of the corresponding triple is not only classical but also it is trivial. In the last section, we define quantum isomorphism between a pair of $2$-graphs. We show that the question whether a pair of $2$-graphs is quantum isomorphic or not is intimately related to the quantum automorphism group of the ‘disjoint union’ of the pair (see Corollary 5.9). The results of the fifth section is a careful ‘extension’ of the results already explored in the literature (see for example [Qinfo]) in the context of usual $1$-graphs. Then we produce a ‘non-trivial’ example of a pair of $2$-graphs that are not quantum isomorphic (see Theorem 5.10). This example of course is completely new. To produce such a pair we consider a $2$-graph with ‘maximal’ quantum symmetry and a $2$-graph with ‘minimal’ quantum symmetry. We end this paper by extending the free wreath product result ([bichon2004free]*Theorem 4.2) for $2$-graphs.
@@ 2.1. Higher rank graphs
Definition 2.2 (see [NYJM]*Definition 1.1).
A $k$-graph (rank $k$ graph or higher rank graph) $(\Lambda,d)$ consists of a countable small category $\Lambda$ (with target and source maps $t$ and $s$ respectively) together with a functor $d:\Lambda\rightarrow\mathbb{N}^{k}$ satisfying the factorization property: for every morphism $\lambda\in\Lambda$ and $m,n\in\mathbb{N}^{k}$ with $d(\lambda)=m+n,$ there are unique morphisms $\mu,\nu\in\Lambda$ such that $\lambda=\mu\nu$ and $d(\mu)=m,\ d(\nu)=n.$
@@
Definition 2.4 (see [NYJM]*Definition 1.9).
Let $f:\mathbb{N}^{l}\rightarrow\mathbb{N}^{k}$ be a monoid morphism. If $(\Lambda,d)$ is a $k$-graph we may form the $l$-graph $f^{\ast}(\Lambda)$ as follows: $f^{\ast}(\Lambda)=\{(\lambda,n):d(\lambda)=f(n)\}$ with $d(\lambda,n)=n,s(\lambda,n)=s(\lambda)\ {\rm and}\ t(\lambda,n)=t(\lambda)$.
Now let us specialize to the case where $k=2$. We shall take an equivalent viewpoint of $2$-graphs which will be amenable to the study of quantum symmetry. This viewpoint is described in [NYJM]*Section 6. Here we shall give the bare bones. Roughly speaking, any $2$-graph can be constructed from two $1$-graphs with common set of vertices and commuting vertex matrices. Again we remind the reader that in our case, since ${\rm obj}(\Lambda)$ is finite, the number of common vertices is finite.
$2$-graphs
Let $\mathcal{G}_{1}=(V,E_{1})$, $\mathcal{G}_{2}=(V,E_{2})$ be two bidirected $1$-graphs with common set of vertices $V$ and without multiple edges, loops or isolated vertices such that the corresponding vertex matrices $\varepsilon_{1},\varepsilon_{2}$ commute. The commutation of $\varepsilon_{1}$ and $\varepsilon_{2}$ guarantees the existence of a bijection $\theta:E_{1}\star E_{2}\rightarrow E_{2}\star E_{1}$ where:
$E_{1}\star E_{2}$ =$=\left\{(e^{1}_{\alpha},e^{2}_{\beta})\in E_{1}\times E_{2}|s(e^{1}_{\alpha})=t(e^{2}_{\beta})\right\},$
$E_{2}\star E_{1}$ =$=\left\{(e^{2}_{\alpha},e^{1}_{\beta})\in E_{2}\times E_{1}|s(e^{2}_{\alpha})=t(e^{1}_{\beta})\right\}.$
Given such a triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$, one can construct a $2$-graph $\Lambda$ in the following way: The set of objects ${\rm obj}(\Lambda)$ or $\Lambda^{0}$ is the set of common vertices of the graphs $\mathcal{G}_{1},\mathcal{G}_{2}$. The morphisms of degree $(1,0)$ and $(0,1)$ are determined by the edge sets $E_{1},E_{2}$. The higher degree morphisms are built step by step using the bijection $\theta$. For details of the construction, we again refer the readers to [NYJM]*Section 6. Conversely, given a $2$-graph $\Lambda$, one can construct a triple. Roughly one takes the set $\Lambda^{(0,0)}$ to be the set of common vertices $V$. Then the morphisms $\Lambda^{(1,0)}$ and $\Lambda^{(0,1)}$ together with the target and source maps defines the constituent $1$-graphs $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$. The bijection $\theta$ is then built out of the morphisms $\Lambda^{(1,1)}$ using the unique factorization property of the morphisms. A triple $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ corresponding to a $2$-graph $\Lambda$ will be called the defining triple of $\Lambda$.
@@
$(1)$ A bijection between the sets $E_{1}\star E_{2}$ and $E_{2}\star E_{1}$ is guaranteed precisely because the vertex matrices commute. Also, $\theta(e^{1}_{\alpha},e^{2}_{\beta})=(e^{2}_{\mu},e^{1}_{\nu})$ implies
$\begin{matrix}s(e^{1}_{\alpha})=t(e^{2}_{\beta});&s(e^{2}_{\mu})=t(e^{1}_{\nu});\\ t(e^{1}_{\alpha})=t(e^{2}_{\mu});&s(e^{2}_{\beta})=s(e^{1}_{\nu}).\end{matrix}$
@@
$(2)$ We also mention that we restrict our attention to the $2$-graphs where the individual $1$-graphs are bidirected without multiple edges, loops or isolated vertices for technical reasons. We assume that the $1$-graphs are without multiple edges and loops to consider ${\rm Aut}^{+}(\mathcal{G}_{1}),{\rm Aut}^{+}(\mathcal{G}_{2})$ [See Lemma 2.18]. The assumption that the graphs are bidirected and without isolated vertices will be used in the proof of Theorem 3.5 [See Section 3] and Lemma 5.3 [See Section 5].
@@
Now let $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ and $(\mathcal{G}_{1}^{\prime},\mathcal{G}^{\prime}_{2},\theta^{\prime})$ be two triples. We denote the common vertex set of the triples by $V,V^{\prime}$ respectively; the vertex matrices of $\mathcal{G}_{l}$ by $\varepsilon_{l}$ and vertex matrices of $\mathcal{G}^{\prime}_{l}$ by $\varepsilon_{l}^{\prime}$ for $l=1,2$ respectively; the edge sets of $\mathcal{G}_{l}$ and $\mathcal{G}_{l}^{\prime}$ by $E_{l}$ and $E_{l}^{\prime}$ for $l=1,2$ respectively.
@@
Two triples $(\mathcal{G}_{1},\mathcal{G}_{2},\theta)$ and $(\mathcal{G}_{1}^{\prime},\mathcal{G}_{2}^{\prime},\theta^{\prime})$ as above are said to be equivalent if the following holds:
  1. (i)
    There is a bijection $T:V\rightarrow V^{\prime}$. Consequently, it induces a linear map $T_{\ast}:C(V^{\prime})\rightarrow C(V)$.
  2. (ii)
    If we denote the permutation matrix of the linear map $T_{\ast}$ again by $T$ with respect to standard bases of $C(V)$ and $C(V^{\prime})$, then $T\varepsilon_{l}^{\prime}=\varepsilon_{l}T$ for $l=1,2$.
  3. (iii)
    Consider the map $T^{\otimes 4}_{\ast}\equiv T_{\ast}\otimes T_{\ast}\otimes T_{\ast}\otimes T_{\ast}:C(V^{\prime}\times V^{\prime}\times V^{\prime}\times V^{\prime})\rightarrow C(V\times V\times V\times V)$. Thanks to condition (ii), $T^{\otimes 4}_{\ast}$ is also a bijection between both $C(E_{1}^{\prime}\star E_{2}^{\prime})$, $C(E_{1}\star E_{2})$ and $C(E_{2}^{\prime}\star E_{1}^{\prime})$, $C(E_{2}\star E_{1})$. We have the following commutative diagram:
[…]
@@
We abbreviate the last condition of the above definition as $\theta$compatibility. The isomorphism classes of $2$-graphs depend only on the equivalence classes of triples as defined above. We have included a brief sketch of the proof. We believe that it must be known. But since we could not find an explicit mention of this result in the literature, we have decided to keep the result in the Appendix (see Theorem A.1).
@@ 2.9. Compact quantum groups and their actions
The term quantum group is quite general. We make it clear that we work with compact quantum groups (CQG in short) ala S.L. Woronowicz ([Woro]). The literature on CQG and their actions on $C^{\ast}$-algebras are aplenty by now. The reader can refer to the articles [DeCom, Soltan] for details on the actions of CQG’s on $C^{\ast}$-algebras. Here we briefly recall what is relevant for us. The following are mainly taken from [edin].
@@
A compact quantum group $\mathbb{G}$ is a pair $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ such that $C(\mathbb{G})$ is a unital $C^{\ast}$-algebra and $\Delta_{\mathbb{G}}:C(\mathbb{G})\rightarrow C(\mathbb{G})\otimes C(\mathbb{G})$ is a unital $C^{\ast}$- homomorphism satisfying
  1. (i)
    $({\rm id}\otimes\Delta_{\mathbb{G}})\circ\Delta_{\mathbb{G}}=(\Delta_{\mathbb{G}}\otimes{\rm id})\circ\Delta_{\mathbb{G}};$
  2. (ii)
    ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(1\otimes C(\mathbb{G}))\}$ and ${\rm Span}\{\Delta_{\mathbb{G}}(C(\mathbb{G}))(C(\mathbb{G})\otimes 1)\}$ are dense in $C(\mathbb{G})\otimes C(\mathbb{G})$.
@@
  1. (i)
    $(\alpha\otimes{\rm id})\circ\alpha=({\rm id}\otimes\Delta_{\mathbb{G}})\circ\alpha$;
  2. (ii)
    Span$\{\alpha({\cal C})(1\otimes\Delta_{\mathbb{G}})\}$ is dense in ${\cal C}\otimes C(\mathbb{G})$;
  3. (iii)
    The $\ast$-algebra generated by the set $\{(\omega\otimes{\rm id})\circ\alpha({\cal C}):\omega\in{\cal C}^{\ast}\}$ is norm-dense in $C(\mathbb{G})$.
Definition 2.13 ([Bichon]*Definition 2.1).
Given a unital $C^{\ast}$-algebra ${\cal C}$, the quantum automorphism group of ${\cal C}$ is a CQG $(C(\mathbb{G}),\Delta_{\mathbb{G}})$ acting faithfully on ${\cal C}$ satisfying the following universal property:
@@
In general, the universal object may fail to exist in the above category. To ensure its existence, one typically assumes that the action preserves some fixed state on the $C^{\ast}$-algebra. However, for certain finite dimensional $C^{\ast}$-algebras, such a state preserving condition is automatic. We will not go into further details here as we will not be using it in this paper. For more details, the reader may consult [Wang].
@@
If we take the space of $n$ points $X_{n}$ then the quantum automorphism group of the $C^{\ast}$-algebra $C(X_{n})$ is given by the CQG (denoted by $S_{n}^{+}$) whose underlying $C^{\ast}$-algebra is the universal $C^{\ast}$ algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ satisfying the following relations (see [Wang]*Theorem 3.1):
@@
Any $n\times n$ matrix $((q_{ij}))_{i,j=1,\ldots,n}$ with $q_{ij}$’s in a $C^{\ast}$-algebra satisfying the above relations is called a magic unitary. The coproduct on the generators is given by $\Delta(q_{ij})=\sum_{k=1}^{n}q_{ik}\otimes q_{kj}$. It is worth mentioning that the CQG $S_{n}^{+}$ is a particular type of CQG known as the compact matrix pseudogroup (see [Woro]). The $\ast$-algebra generated by $\{q_{ij}\}_{i,j=1,\ldots,n}$ is the canonical dense Hopf $\ast$-algebra and the antipode $\kappa$ is given on the generators by $\kappa(q_{ij})=q_{ji}$.